### v2

# Compressing Language Models for Specialized Domains

Abstract
Language models (LMs) excel at tasks across diverse domains, yet require substantial computational resources during inference. Compression techniques such as pruning and quantization offer a practical path towards efficient LM deployment, exemplified by their ability to preserve performance on general-purpose benchmarks. However, general-purpose LM compression methods can negatively affect performance in specialized domains (e.g. biomedical or legal). Recent work has sought to address this issue, but requires a computationally expensive full-parameter fine-tuning pipeline. To this end, we propose MixCal, a novel calibration method designed to improve the in-domain performance of compressed LMs in a post-training setting. Through extensive experimentation, we demonstrate that MixCal substantially outperforms existing approaches on domain-specific tasks and preserves general performance. Notably, these performance gains are achieved while also reducing the computational cost of LM compression. [^1] https://github.com/mlsw/domain-compression
Compressing Language Models for Specialized Domains
Miles Williams♢♠   George Chrysostomou   Vitor Jeronymo   Nikolaos Aletras University of Sheffield Enterprise AI Services, AstraZeneca {mwilliams15, n.aletras}@sheffield.ac.uk
1 Introduction
Language models (LMs) have demonstrated remarkable performance across tasks from a range of domains (Walsh et al., 2025; Guo et al., 2025; Kamath et al., 2025). Behind this success lies a recipe with two key ingredients: highly parameterized models and extensive training. However, the vast scale of these models presents substantial challenges in their deployment and application (Treviso et al., 2023; Zhu et al., 2024). Luccioni et al. (2024) suggest that the trend towards general-purpose models has introduced substantial yet potentially unnecessary inference costs.
Model compression techniques, such as quantization and pruning, are foundational approaches aimed at reducing the computational footprint of LMs during inference (Zhu et al., 2024). Quantization represents weights (and/or activations) with lower precision, while pruning removes less important weights. Notably, recent work has shown the successful application of quantization (Frantar et al., 2023; Lin et al., 2024) and pruning (Frantar and Alistarh, 2023; Sun et al., 2024) to general-purpose LMs without any additional training.
Figure 1: MixCal (§3) is applied within the calibration phase of model compression, leveraging a combination of activations from generic and domain-specific data.
LM compression studies typically focus on preserving general-purpose performance, i.e. language modeling and commonsense reasoning capabilities (Frantar and Alistarh, 2023; Ma et al., 2023; Sun et al., 2024). However, in practice, LMs may be deployed within only one particular domain, e.g. biomedical or legal (Labrak et al., 2024; Colombo et al., 2024; Ling et al., 2024; Chen et al., 2024). This scenario unlocks new paths towards improving inference efficiency by extracting domain-specific LMs from general-purpose models.
Zhang et al. (2024) proposed D-Pruner, a pruning method aiming to preserve weights that are influential to both domain-specific and general capabilities. To identify such weights, D-Pruner leverages the gradient information from a composite loss function that incorporates general weight importance scores. However, this requires a full-parameter fine-tuning pipeline for the LM, thus incurring substantial computational costs.
The majority of post-training LM compression methods rely upon calibration data, a small amount of data used to aid the analysis of layer activations. Recent work has shown that calibration data can impact task performance (Williams and Aletras, 2024; Bandari et al., 2024; Williams et al., 2025). Inspired by this line of work, we investigate how to effectively leverage domain-specific data for calibration, aiming to maximize in-domain performance without sacrificing general capabilities. Our main contributions are as follows:
  1. 1.
    We find that in-domain calibration data can play an important role in LM pruning, maximizing performance retention on domain-specific tasks.
  2. 2.
    We propose MixCal, a novel calibration method for Hessian-based compression, enabling the effective use of in-domain calibration data (Figure 1). Our approach outperforms existing general and domain-specific pruning methods while offering greater computational efficiency.
2 Related Work
Quantization.
The objective of quantization is to represent weights (and optionally activations) using fewer bits. This reduction in precision reduces memory requirements, and typically enables inference speedups (Gholami et al., 2021). Beyond their scale, contemporary LMs pose unique challenges for effective quantization, including the existence of high-magnitude outlier features (Bondarenko et al., 2021; Dettmers et al., 2022). Recent directions include: holding outlier weights in higher precision (Dettmers et al., 2022), Hessian-based weight sensitivity (Frantar et al., 2023), searching for optimal clipping thresholds (Wei et al., 2023), or combinations of these approaches (Lin et al., 2024; Dettmers et al., 2024).
Pruning.
The aim of neural network pruning is to remove less important weights, therefore reducing the overall model size (LeCun et al., 1989). Pruning can be performed at the level of individual weights (unstructured), within groups of weights (semi-structured), or entire dimensions (structured) (Han et al., 2015; Mishra et al., 2021; Ma et al., 2023). In particular, 2:4 semi-structured sparsity (i.e. pruning two weights in every block of four) enables enhanced inference performance on NVIDIA GPUs (Mishra et al., 2021). However, the extensive size of Transformer-based (Vaswani et al., 2017) LMs presents challenges in pruning them optimally (Hassibi et al., 1993). Recent work has instead decomposed LM pruning into a sequential layer-wise approach, demonstrating remarkable performance retention, even at high sparsity levels (Frantar and Alistarh, 2023; Sun et al., 2024; Yin et al., 2024).
Domain-specific pruning.
Early work focused on pruning deep neural networks for specific tasks (Han et al., 2015; Molchanov et al., 2017). This trend continued (Sanh et al., 2020; Lagunas et al., 2021; Kwon et al., 2022) following the advent of BERT (Devlin et al., 2019). However, the shift towards general-purpose LMs (Brown et al., 2020; Dredze et al., 2024) has led to a focus on preserving general performance (Frantar and Alistarh, 2023; Ma et al., 2023; Sun et al., 2024), i.e. language modeling and reasoning. Most recently, Zhang et al. (2024) proposed D-Pruner for domain-specific pruning, leveraging general weight importance to form a domain-specific training loss. However, this requires an expensive full-parameter fine-tuning pipeline, yet does not consistently outperform general-purpose pruning methods.
Calibration data.
In a post-training setting, model compression usually relies upon calibration data (Wan et al., 2024). Calibration data consists of a small number of unlabeled examples for the generation of layer activations (Nagel et al., 2020; Hubara et al., 2021). Typically, these examples are randomly sampled from web text or pre-training datasets (e.g. C4; Raffel et al., 2020). Recent work has illustrated the influential role that calibration data can play, impacting the downstream performance of compressed models (Williams and Aletras, 2024; Williams et al., 2025). However, Bandari et al. (2024) suggest that pruning using downstream task data does not necessarily outperform generic data on the corresponding task.
3 MixCal
3.1 Preliminaries
The Optimal Brain Surgeon (OBS; Hassibi et al., 1993) algorithm leverages second-order derivatives to accurately prune weights from a neural network. These second-order derivatives, which indicate the curvature of the loss function with respect to the weights, are organized in a square matrix known as the Hessian $\mathbf{H}$. Based on the Hessian, the OBS algorithm iteratively removes the weight $w_{m}$ with the lowest saliency $\varepsilon_{m}$, followed by applying the optimal update for the remaining weights $\boldsymbol{\delta}_{m}$:
$\varepsilon_{m}=\frac{1}{2}\frac{w^{2}_{m}}{[\mathbf{H}^{-1}]_{mm}},\quad\boldsymbol{\delta}_{m}=-\frac{w_{m}}{[\mathbf{H}^{-1}]_{mm}}\cdot\mathbf{H}^{-1}_{:,m}$
Frantar and Alistarh (2022) reformulate pruning as a case of the layer-wise compression problem, while retaining the OBS weight update procedure. Given a layer with input activations $\mathbf{X}$, the aim is to minimize the error between the original layer weights $\mathbf{W}$ and newly compressed weights $\mathbf{\widehat{W}}$:
$\underset{\mathbf{\widehat{W}}}{\arg\min}\;\|\mathbf{X}\mathbf{W}-\mathbf{X}\mathbf{\widehat{W}}\|^{2}_{F}$
As the layer input activations are derived from a fixed set of calibration data, the layer outputs (i.e. $\mathbf{Y}=\mathbf{XW}$) are also fixed. Consequently, the layer-wise Hessian is computed as:
$\mathbf{H}=2\mathbf{X}^{\top}\mathbf{X}$
3.2 Motivation
Language models are trained over diverse and expansive corpora. Accordingly, a random sample of generic calibration data often proves sufficient to approximate the Hessian (Frantar and Alistarh, 2023). These generic samples can also help preserve general capabilities (Bandari et al., 2024). However, LM weights additionally encode domain-specific knowledge (Singhal et al., 2023). For example, Figure 2 illustrates that a feature may be sensitive in a target domain yet appear unimportant under generic data. Therefore, we hypothesize that compressing LMs for specialized domains while maintaining general performance requires preserving weights that are important under both settings. To this end, we propose MixCal (Algorithm 1).
3.3 Mixing Activations
First, we extend layer-wise compression to a multi-objective setting. The standard layer-wise reconstruction loss for a layer with weights $\mathbf{W}$, compressed weights $\widehat{\mathbf{W}}$, and input activations $\mathbf{X}$ is:
$\mathcal{L}(\widehat{\mathbf{W}},\mathbf{X})=\|\mathbf{X}\mathbf{W}-\mathbf{X}\widehat{\mathbf{W}}\|^{2}_{F}$
We introduce two distinct sources of calibration data, a domain-specific dataset $\mathcal{D}_{d}$, and a general-purpose dataset $\mathcal{D}_{g}$. Samples from these datasets are used to form the domain-specific and general-purpose activations, $\mathbf{X}_{d}$ and $\mathbf{X}_{g}$, respectively. We then define a combined loss, which balances reconstruction between both calibration datasets:
$\mathcal{L}_{\text{weighted}}=\alpha\mathcal{L}(\widehat{\mathbf{W}},\mathbf{X}_{d})+\beta\mathcal{L}(\widehat{\mathbf{W}},\mathbf{X}_{g})$
To control the relative contribution from each loss term, we introduce the weight coefficients $\alpha$ and $\beta$. We impose $\alpha+\beta=1$ to avoid an arbitrary scaling, setting $\beta=1-\alpha$. The loss is therefore parameterized by a single coefficient, $\alpha\in[0,1]$. This loss function yields the following Hessian:
$\mathbf{H}_{\text{weighted}}=2(\alpha\mathbf{X}_{d}^{\top}\mathbf{X}_{d}+\beta\mathbf{X}_{g}^{\top}\mathbf{X}_{g})$
However, this formulation treats the domain-specific and general-purpose activations independently. The Hessian $\mathbf{H}_{\text{weighted}}$ is a weighted sum of the domain-specific and general-purpose Hessians, namely $2\mathbf{X}_{d}^{\top}\mathbf{X}_{d}$ and $2\mathbf{X}_{g}^{\top}\mathbf{X}_{g}$. These terms only capture the second-order structure for their corresponding dataset. Consequently, $\mathbf{H}_{\text{weighted}}$ does not encode interactions between the two datasets.
Figure 2: The Hessian at layer 16 of Mistral NeMo 12B, computed with (a) generic calibration data, and (b) domain-specific calibration data. For clarity, we present the magnitude of the elements for the first 32 features.
We therefore propose a strategy which also incorporates such cross-dataset interactions. Concretely, we construct a mixed activation matrix $\mathbf{Z}$ that combines the activations from each of the calibration datasets. This is used to form the loss $\mathcal{L}_{\text{mix}}$:
$\mathbf{Z}=\sqrt{\alpha}\mathbf{X}_{d}+\sqrt{\beta}\mathbf{X}_{g},\quad\mathcal{L}_{\text{mix}}=\mathcal{L}(\widehat{\mathbf{W}},\mathbf{Z})$
Expanding the Hessian for $\mathcal{L}_{\text{mix}}$ illustrates the additional cross term that is introduced:
$\mathbf{H}_{\text{mix}}$ $=\mathbf{H}_{\text{weighted}}+2\sqrt{\alpha\beta}(\mathbf{X}_{d}^{\top}\mathbf{X}_{g}+\mathbf{X}_{g}^{\top}\mathbf{X}_{d})$
The contribution from this term can be interpreted as how strongly the domain-specific and general-purpose activations align.
Finally, we note a connection to the training-time data augmentation concept of mixup (Zhang et al., 2018), which forms linear combinations of pairs of examples and their labels. Mixup has previously been applied directly to activations (Verma et al., 2019) and Transformer LMs (Sun et al., 2020). While our approach also forms linear combinations of activations, we use unlabeled examples solely to approximate the Hessian for compression.
3.4 Implementation
In practice, we compute the Hessian using the form $\mathbf{H}_{\text{mix}}=2\mathbf{Z}^{\top}\mathbf{Z}$. When the total number of calibration examples is held constant, this can reduce the cost of calibration, since each update to the Hessian incorporates two examples simultaneously. We analyze the resulting efficiency gains in §5.3.
To empirically validate our approach, we integrate MixCal with SparseGPT (Frantar and Alistarh, 2023), a popular Hessian-based pruning algorithm. However, we emphasize that MixCal is not tied to any specific compression algorithm (see Figure 7, Appendix A.1 for results with GPTQ-M).
4 Experimental Setup
4.1 Compression Methods
Pruning methods can be formulated as a function that computes a saliency score $\mathbf{S}_{ij}$ for each weight $\mathbf{W}_{ij}$ in a given layer. They optionally use the layer input activations $\mathbf{X}$, derived from calibration data. We adopt the following methods as baselines.
Magnitude (Janowsky, 1989; Han et al., 2015).
Based on the assumption that removing the smallest weights will have the least effect, magnitude pruning simply uses the weight magnitude for saliency:
$\mathbf{S}_{ij}=|\mathbf{W}_{ij}|$
SparseGPT (Frantar and Alistarh, 2023).
Building upon the OBS procedure, SparseGPT offers an efficient iterative approximation. The saliency metric is computed as follows, where $\lambda$ is a dampening factor to enable inversion of the Hessian:
$\mathbf{S}_{ij}=\left[|\mathbf{W}|^{2}/\text{diag}\left((\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\right)\right]_{ij}$
Wanda (Sun et al., 2024).
Improving upon the computational efficiency of SparseGPT, the Wanda pruning metric approximates the diagonal of the inverse Hessian via the $\ell_{2}$ norm of the activations:
$\mathbf{S}_{ij}=|\mathbf{W}_{ij}|\cdot\|\mathbf{X}\|_{2}$
D-Pruner (Zhang et al., 2024).
D-Pruner is a domain-specific pruning method. The first step of D-Pruner is to compute the general importance $\mathbf{G}_{ij}$ using a general dataset $\mathcal{D}_{g}$, similar to SparseGPT. Second, it uses a composite loss function $\mathcal{L}_{\text{DP}}$ to identify weights that are important for both general and domain-specific knowledge. This consists of the cross-entropy loss $\mathcal{L}_{\text{CE}}$ with a regularization term controlled by hyperparameter $\lambda_{g}$, where $\mathbf{W}^{\prime}$ is the updated weight matrix:
$\mathcal{L}_{\text{DP}}\approx\mathcal{L}_{\text{CE}}+\lambda_{g}\sum_{i,j}\mathbf{G}_{ij}(\mathbf{W}^{\prime}_{ij}-\mathbf{W}_{ij})^{2}$
The gradients are computed using a full-parameter fine-tuning pipeline (Lv et al., 2024). Given a domain-specific dataset $\mathcal{D}_{d}$, the saliency is:
$\mathbf{S}_{ij}\approx\left|\frac{\partial\mathcal{L}_{\text{DP}}(\mathcal{D}_{d})}{\partial\mathbf{W}_{ij}}\mathbf{W}_{ij}+\frac{1}{2}\left[\frac{\partial\mathcal{L}_{\text{DP}}(\mathcal{D}_{d})}{\partial\mathbf{W}_{ij}}\mathbf{W}_{ij}\right]^{2}\right|$
Algorithm 1 MixCal (simplified).
1:Domain dataset $\mathcal{D}_{d}$, generic dataset $\mathcal{D}_{g}$, weight $\alpha\in[0,1]$, embedding matrix $\mathbf{U}$, and weight matrices $\mathcal{W}$.
2:for $i\leftarrow 1$ to $|\mathcal{D}_{d}|$ do $\triangleright$ Since $|\mathcal{D}_{d}|=|\mathcal{D}_{g}|$.
3:  $(\mathbf{X}_{d,i},\,\mathbf{X}_{g,i})\leftarrow(\mathcal{D}_{d,i}\cdot\mathbf{U},\,\mathcal{D}_{g,i}\cdot\mathbf{U})$
4:end for
5:for $\mathbf{W}\in\mathcal{W}$ do
6:  $\mathbf{H}\leftarrow\mathbf{0}$
7:  for $i\leftarrow 1$ to $|\mathcal{D}_{d}|$ do
8:   $\mathbf{Z}_{i}\leftarrow\sqrt{\alpha}\mathbf{X}_{d,i}+\sqrt{1-\alpha}\mathbf{X}_{g,i}$
9:   $\mathbf{H}\leftarrow\frac{i-1}{i}\mathbf{H}+\frac{2}{i}(\mathbf{Z}_{i}^{\top}\mathbf{Z}_{i})$
10:  end for
11:  $\mathbf{W}\leftarrow\text{Compress}(\mathbf{W},\,\mathbf{H})$ $\triangleright$ E.g. via SparseGPT.
12:  for $i\leftarrow 1$ to $|\mathcal{D}_{d}|$ do
13:   $(\mathbf{X}_{d,i},\,\mathbf{X}_{g,i})\leftarrow(\mathbf{X}_{d,i}\mathbf{W},\,\mathbf{X}_{g,i}\mathbf{W})$
14:  end for
15:end for
Figure 3: The average benchmark accuracy when pruning to 50% sparsity, relative to the original model.
GPTQ-M (Frantar et al., 2025).
GPTQ adopts a Hessian-based weight sensitivity metric for quantization (Frantar et al., 2023). We select this method to enable a fair comparison with SparseGPT, which uses GPTQ for joint sparsification and quantization. We include improvements to the original algorithm suggested by Frantar et al. (2025). Specifically, this identifies optimal group-wise clipping thresholds, similar to AWQ (Lin et al., 2024). For clarity, we refer to this improved method as GPTQ-M.
Compression configurations.
Guided by prior work (Sun et al., 2024; Frantar et al., 2025), we focus our experiments on the following settings:
  • 50% (unstructured) sparsity. First, we experiment with individually pruning half of all layer weights, offering the highest possible granularity.
  • 2:4 (semi-structured) sparsity. We then examine pruning at the granularity of two weights in every group of four, enabling enhanced GPU inference performance (Mishra et al., 2021).
[…]
4.2 Domains and Tasks
To assess the efficacy of our approach on downstream tasks, we experiment with two of the most widely explored domains in NLP, the biomedical (Lee et al., 2019; Gu et al., 2021; Luo et al., 2022; Singhal et al., 2023, 2025) and legal (Chalkidis et al., 2019; Zheng et al., 2021; Henderson et al., 2022; T.y.s.s et al., 2024; Niklaus et al., 2024) domains. See Appendix C for concrete task examples.
Biomedical.
We use the MultiMedQA benchmark (Singhal et al., 2023), specifically the PubMedQA (Jin et al., 2019), MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022) tasks, and relevant subsets from MMLU (Hendrycks et al., 2021) (anatomy, clinical knowledge, college medicine, medical genetics, professional medicine, college biology). To assess language modeling performance, we use the BioLaySumm PLOS dataset (Goldsack et al., 2022, 2023) comprising biomedical articles.
Legal.
We follow Colombo et al. (2024) in using Legal-MMLU, covering jurisprudence, professional law, and international law specialties (Hendrycks et al., 2021). We also use the CaseHOLD (Zheng et al., 2021) and ECtHR (Task A) (Chalkidis et al., 2019) datasets from the LexGLUE benchmark (Chalkidis et al., 2022), comprising US Supreme Court opinions and European Court of Human Rights cases, respectively. To evaluate language modeling performance, we use the BillSum dataset (Kornilova and Eidelman, 2019) of US Congressional and California state bills.
General.
To assess general performance, we use all commonsense reasoning tasks adopted by Frantar and Alistarh (2023) and Sun et al. (2024): ARC (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), LAMBADA (Paperno et al., 2016), OpenBookQA (Banerjee et al., 2019), PIQA (Bisk et al., 2020), RTE (Dagan et al., 2006), StoryCloze (Mostafazadeh et al., 2016), and WinoGrande (Sakaguchi et al., 2021). We use WikiText-2 (Merity et al., 2017) to assess language modeling.
4.3 Calibration Data
Data sources.
To create our general-purpose calibration sets, we follow Dettmers et al. (2024) in using RedPajama (Weber et al., 2024), an open reproduction of the LLaMA training data. For the domain-specific calibration sets, we use MultiMedQA (biomedical) and LexGLUE (legal). In all cases, we sample data from the training splits only.
Data quantity.
For a fair comparison between compression methods, we use 1024 calibration examples. As D-Pruner consists of two distinct stages to identify general and domain-specific weight importance, we allow 1024 examples from each dataset to better match the original work (Zhang et al., 2024). For our own method, we simply select half of the examples (i.e. 512) from each dataset.
Sampling.
We randomly sample segments of 2048 tokens following Frantar et al. (2023), avoiding any selection bias. In the case of the domain-specific datasets, which may contain shorter examples, we follow Touvron et al. (2023) in concatenating examples for a consistent length. We repeat the sampling process to create five distinct calibration sets, used to assess the variance in performance.
4.4 Models
We experiment with popular open-weights LMs, covering different model families and sizes: (1) Llama 3.2 3B and 3.1 8B (Grattafiori et al., 2024), (2) Gemma 2 2B and 9B (Riviere et al., 2024), and (3) Mistral NeMo 12B (2407) (Jiang et al., 2024).
5 Results & Discussion
5.1 Pruning
Figure 3 presents the benchmark accuracy when pruning to 50% sparsity, relative to the original model. [^2] The D-Pruner (Zhang et al., 2024) implementation only supports models with the Llama architecture. See Limitations. We report the mean value and standard deviation across five calibration sets. For brevity, we present the average performance across domain-specific models. We additionally present complete results across all models in Appendix E.
A note on hyperparameters.
To maximize the performance of the D-Pruner baseline, we perform an extensive hyperparameter search across $\lambda_{g}\in\{0.1,0.01,0.001\}$ and group size $\in\{\text{None},128\}$ for each model and domain. We then present results for only the best performing combinations. We present complete results across all hyperparameters in Appendix E. In contrast, we do not optimize the hyperparameter for our approach ($\alpha$) and simply use 0.5 across all models and domains. We ablate the impact of this hyperparameter in Appendix A.3.
MixCal benefits domain performance.
We observe that across both biomedical and legal domains, our approach consistently outperforms all other compression methods. For example, we observe that MixCal achieves an average relative accuracy of 88.6% on the legal benchmark for Llama 3.1 8B. In comparison, SparseGPT and D-Pruner see 72.9% and 77.2%, respectively. For the biomedical benchmark, a similar trend can be observed. MixCal achieves 86.5%, while SparseGPT and D-Pruner reach 75.3% and 81.0%, respectively. This indicates that MixCal can be effective at identifying weights influential to in-domain performance.
Figure 4: Average accuracy when applying 4-bit quantization and 2:4 sparsity, relative to the original model.
General performance is comparable in all cases.
In addition to substantial improvements on domain-specific benchmarks, general performance remains similar to the general-purpose pruning baselines. With Llama 3.1 8B, MixCal achieves 92.6% relative accuracy for general tasks on average. In comparison, SparseGPT achieves a slightly lower value of 92.2%, while D-Pruner reaches only 85.7%. Intriguingly, we note that MixCal achieves higher performance than even the general-purpose methods for most model families. For example, MixCal sees an average of 94.5% with Gemma 2 9B, while SparseGPT reaches 93.7%. This suggests that MixCal can effectively establish domain-specific features without sacrificing general performance.
Performance gains are generally model-agnostic.
Finally, we observe that the performance benefits of MixCal are similar irrespective of the model size and family. For example, we consider two similarly sized models, Llama 3.1 8B and Gemma 2 9B. In the biomedical benchmark, we observe an 11.2 and 7.5 point increase in relative accuracy over SparseGPT, respectively. For the legal benchmark, we see a 15.7 and 16.1 point increase. This suggests that the performance gains from MixCal are independent of the model family.
Language modeling follows a similar trend.
Table 7 (Appendix E) presents perplexity results. Similar to the downstream task experiments, we note that MixCal achieves the best performance on in-domain language modeling. Considering Llama 3.1 8B, MixCal achieves a perplexity of 6.0 compared to 6.7 with D-Pruner for the legal domain. For the biomedical domain, MixCal has a perplexity of 10.7, compared to 14.7 from D-Pruner. The datasets used to evaluate perplexity are not used for calibration, suggesting that MixCal assists with identifying domain-specific features.
MixCal appears to generalize beyond specific tasks.
Across all models and domains, we observe that the performance benefits of MixCal continue to tasks not included in the calibration data. In Table 10 (Appendix E), we present complete per-task results. We consider the MMLU tasks, which do not have training data, and are therefore not represented in the calibration data. For Llama 3.1 8B, MixCal achieves an absolute increase in accuracy of 7.7 points over SparseGPT in the biomedical domain. In the legal domain, MixCal achieves an increase of 2.7 points in accuracy. This suggests that MixCal can identify features that are relevant to the domain, rather than only a specific task.
5.2 Joint Pruning and Quantization
We further examine the performance of our approach when jointly applying pruning and quantization by reusing the same inverse Hessian (Frantar and Alistarh, 2023). This has the advantage of allowing pruning and quantization decisions to influence each other, and enables quantization at almost no extra cost. Figure 4 presents benchmark accuracy when jointly applying 2:4 sparsity with 4-bit quantization, relative to the original model.
MixCal improves in-domain performance while sustaining general performance.
MixCal achieves substantially greater domain performance than SparseGPT for 2:4 sparsity. For example, it sees a relative accuracy of 77.8% on the biomedical benchmark for Gemma 2 9B, versus 58.2% with SparseGPT. For general performance, we observe a similar trend to the pruning results, with MixCal also performing comparably. This illustrates that MixCal can be reliably used with quantization.
5.3 Compression Efficiency
Figure 5: The average duration and peak memory allocated when pruning Llama 3.1 8B with each method, as measured using an NVIDIA A100 80GB GPU.
MixCal does not sacrifice efficiency.
Figure 5 presents the time and memory requirements of each method, which are often limiting factors in practice. First, we observe that MixCal (0.5 hours) reduces the duration of compression compared to SparseGPT (0.8 hours). This is considerably faster than D-Pruner, which on average takes 9.3 hours, 18 times longer than our approach. We also observe that MixCal does not increase the memory required for compression over SparseGPT, with both using up to 20 GB of memory. In contrast, D-Pruner uses up to 65.6 GB of memory, over three times more than our approach. This suggests that MixCal is more practical than D-Pruner, offering lower computational resource requirements.
6 Analysis
6.1 Ablating MixCal
To better understand the role that in-domain data plays in pruning, we conduct an ablation study of MixCal. Table 1 presents the in-domain and general language modeling performance (perplexity) versus a SparseGPT baseline with generic calibration data. We specifically show the effect of (1) adding in-domain data to the calibration data mixture, and (2) additionally using MixCal. In both cases, we use the same equal mixture of in-domain and generic data. Results are averaged across models, with granular results in Table 11, Appendix E.
Method In-domain General Average
Biomedical
SparseGPT 16.4 14.0 15.2
+ In-domain data 15.1 13.8 14.4
+ MixCal (Ours) 14.4 13.4 13.9
Legal
SparseGPT 7.2 14.0 10.6
+ In-domain data 6.7 13.7 10.2
+ MixCal (Ours) 6.7 13.5 10.1
Table 1: Perplexity when pruning to 50% unstructured sparsity with (1) in-domain data, and (2) MixCal, averaged across all models.
Domain-specific data benefits language modeling performance.
We first observe that using a mix of domain-specific and generic calibration data can substantially improve in-domain performance. For example, when using a mix of in-domain data, we find lower perplexity scores for both domains. Specifically, we observe an average perplexity of 15.1 versus 16.4 for the biomedical domain, and 6.7 versus 7.2 for the legal domain. This highlights the influential role played by the calibration data, corroborating Williams and Aletras (2024).
MixCal maximizes overall performance.
Finally, we observe that by applying MixCal, language modeling performance is maximized compared to introducing domain-specific calibration data alone. For example, MixCal achieves a reduction of 0.7 (average) perplexity for the biomedical domain (14.4 versus 15.1). For the legal domain, we note that the average perplexity is equivalent. Considering overall performance, i.e. the average of domain-specific and general performance, MixCal yields the best results. This suggests that the addition of MixCal can balance domain-specific and general performance effectively.
6.2 Performance in Other Languages
To explore whether MixCal generalizes beyond English, we further experiment with a Chinese-language model and benchmarks. We select the Chinese language as it is morphologically distinct from English and well-resourced in terms of models and domain-specific evaluation tasks.
Experimental setup.
We select the Yi 1.5 6B model (Young et al., 2025) as it (1) uses the Llama architecture, enabling experiments with D-Pruner, and (2) achieves strong performance on standard benchmarks. To assess biomedical performance, we use the Comprehensive Medical Benchmark (CMB; Wang et al., 2024). For legal performance, we use the Chinese AI and Law (CAIL2018) challenge dataset (Xiao et al., 2018). Similar to our experimental setup in English, we sample in-domain calibration data from the training split of each benchmark. For general-purpose data, we follow Kurz et al. (2026) in using mC4 (Xue et al., 2021). We use a mixture of Chinese and English to reflect the model pre-training data (Young et al., 2025).
Figure 6: The mean accuracy for Yi 1.5 6B on Chinese-language benchmarks, relative to the original model.
MixCal appears language-agnostic.
Figure 6 presents the relative benchmark accuracy when pruning to 50% sparsity, similar to earlier experiments (§5.1). The results indicate that MixCal outperforms other pruning approaches for domain performance, while maintaining comparable general performance to SparseGPT. For example, MixCal achieves a relative accuracy of 85.6% in the biomedical domain, compared to 70.0% and 79.6% from SparseGPT and D-Pruner, respectively. In the legal domain, the performance retention from MixCal (90.1%) is substantially greater than both SparseGPT (47.1%) and D-Pruner (76.2%). These findings are in line with the English language tasks, suggesting that MixCal is language-agnostic.
7 Conclusion
In this paper, we proposed MixCal as a solution for creating compressed LMs for specialized domains. We empirically validated MixCal using a plethora of pre-trained models and evaluation tasks. Our approach represents a substantial advancement over earlier work such as D-Pruner, offering consistent performance improvements with a smaller computational footprint. We hope that our study will inspire further work towards the efficient deployment of LMs in specialized domains. As future work, we are interested in exploring how synthetic calibration data could be incorporated to further enhance the performance of domain-specific LM compression (Williams et al., 2025).
Limitations
Model selection in the D-Pruner experiments.
The D-Pruner (Zhang et al., 2024) implementation supports only the Llama model architecture. [^3] https://github.com/psunlpgroup/D-Pruner Consequently, we are limited to offering comparisons for only the models using the Llama architecture (i.e. Llama 3.2 3B, Yi 1.5 6B, and Llama 3.1 8B). We emphasize that our approach substantially outperforms D-Pruner across all tested models and domains. Therefore, we expect that this trend would continue for the models not supported by D-Pruner.
Ethical Considerations
Our work enables the efficient and effective compression of LMs for specialized domains. We note that this poses dual-use concerns, as it may enable misuse at a lower cost (Weidinger et al., 2022). However, we emphasize that our approach is unlikely to enhance or introduce new harmful abilities, as the performance of compressed models is constrained by the capabilities of the original.
Acknowledgments
We would like to thank Huiyin Xue and the anonymous reviewers for their invaluable feedback. Additionally, we are sincerely grateful to Huiyin Xue for assistance with the Chinese-language tasks. MW is supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation grant EP/S023062/1. NA is supported by EPSRC grant EP/Y009800/1, part of the RAI UK Keystone projects.
Appendix A Additional Ablation Studies
A.1 4-bit Quantization
MixCal generalizes to quantization.
We further examine the performance and transferability of our approach under the setting of quantization. Figure 7 presents the benchmark accuracy when applying 4-bit quantization with GPTQ-M and MixCal, relative to the original model. This uncovers the extent to which MixCal benefits each method independently in the joint sparsification and quantization experiments (§5.2). We observe that general performance is consistently preserved, with GPTQ-M + MixCal performing comparably to GPTQ-M across all models. This suggests that MixCal is transferable across settings in Hessian-based compression.
MixCal improves in-domain performance.
We observe that MixCal can achieve greater domain performance than GPTQ-M alone. For example, MixCal sees a relative accuracy of 96.6% on the biomedical benchmark for Llama 3.1 8B, versus 94.7% with GPTQ-M (Figure 7). This indicates that MixCal can be reliably used for LM quantization. We note that the performance gains for quantization are smaller than in the pruning experiments. This is expected, as quantization is less sensitive to calibration data at the tested sparsity level and bit width (Williams and Aletras, 2024).
A.2 Performance Across Sparsity Levels
In Figure 8, we examine how MixCal performs compared to the other pruning methods across different sparsity levels. We report the average in-domain and general benchmark performance, with standard deviation denoted by the shaded regions. We first observe that in-domain performance with MixCal consistently remains higher than other approaches, even beyond 50% sparsity. For example, accuracy on the legal benchmark at 60% sparsity is 43.3% with MixCal, compared to 27.3% with SparseGPT. Similarly, general benchmark performance remains comparable to SparseGPT across all sparsity levels. This suggests that MixCal can be used to effectively isolate domain-specific features without sacrificing generic performance.
A.3 MixCal Hyperparameter Analysis
Figure 7: Average accuracy when applying 4-bit quantization, relative to the original model.
Method Hyperparameter Value
D-Pruner Loss Regularization {0.001, 0.01, 0.1}
Group Size {None, 128}
Learning Rate 0.03
GPTQ-M Bits per Weight 4
Dampening 0.01
Group Size 128
Symmetric Quantization Yes
SparseGPT Dampening 0.01
Group Size 128
Sparsity {0.5, 2:4}
Wanda Group Size 1
Sparsity {0.5, 2:4}
Table 2: Hyperparameters used for all compression methods evaluated in our experiments.
In Figure 9, we examine how different contributions from domain-specific and general-purpose data impact benchmark performance, i.e. by varying $\alpha$. Following the experiments in §5.1, we examine this at 50% unstructured sparsity.
Figure 8: Benchmark accuracy when pruning Llama 3.1 8B to different sparsity levels with each method. Standard deviation is denoted by the shaded regions.
Figure 9: Benchmark accuracy relative to the original model when varying the MixCal $\alpha$ hyperparameter for pruning. Standard deviation is denoted by the shaded regions.
Increasing $\boldsymbol{\alpha}$ generally increases in-domain performance.
First, we observe that in-domain performance (i.e. biomedical or legal) increases following a larger contribution from the domain-specific Hessian. Considering Gemma 2 9B for the biomedical domain, we observe that $\alpha=0.1$ achieves an average accuracy of 89.1%, whilst $\alpha=0.9$ sees 93.1%. Notably, we observe that even a small contribution from the in-domain Hessian ($\alpha=0.1$) is enough to offer improvements for in-domain performance. For example, in the legal domain Llama 3.2 3B achieves an average accuracy of 63.1% with $\alpha=0$, yet 71.0% with $\alpha=0.1$.
Maximizing $\alpha$ can harm reasoning performance.
We observe that using $\alpha=1$ can lead to a drop in reasoning performance. For example, when targeting the legal domain, Llama 3.2 3B achieves an average accuracy of 91.5% with $\alpha=0.9$, yet sees 89.3% when $\alpha=1$. This suggests that the use of generic data can benefit reasoning performance.
Appendix B Hyperparameters
Table 2 presents the hyperparameters of all the compression methods in our experiments. In general, we adopt the optimal hyperparameters used for each method in the original work. For completeness, we also present results for every tested D-Pruner hyperparameter combination in Table 9.
Appendix C Data & Processing
Table 3 shows the splits for the datasets used in our experiments, split by category. We also show examples of how the domain-specific tasks are formatted for zero-shot LM evaluation. These are presented in Tables 4 and 5 for the legal and biomedical domains, respectively.
In the majority of cases, we use the task datasets exactly as implemented by the EleutherAI LM Evaluation Harness (Gao et al., 2024). We highlight the exceptions where additional preprocessing was required, below:
  • CaseHOLD and ECtHR (Task A). We adopt the versions of these datasets as provided by the LexGLUE benchmark (Chalkidis et al., 2022). To enable evaluation in a zero-shot setting, we adapt the prompts from Chalkidis (2023).
  • ECtHR (Task A), CAIL2018, and CMB. We additionally filter examples with multiple labels from these datasets, following prior work towards adapting existing tasks for few-shot LM evaluation (Guha et al., 2023).
  • CAIL2018 and CMB. For Chinese-language evaluation tasks, we use the prompts presented in Table 6, Appendix C.
  • BillSum and BioLaySumm PLOS. Due to the extensive size of the test split in these datasets, we follow Frantar et al. (2023) in using the first 256 examples to assess perplexity. We highlight that perplexity is a stable metric which can be assessed using only a small number of examples (Dettmers and Zettlemoyer, 2023).
Name Train Val. Test
Biomedical
BioLaySumm PLOS 24,773 1,376 256
CMB (Chinese) 231,902 228 9,325
PubMedQA 450 50 500
MedMCQA 182,822 4,183 4,183
MedQA 10,178 1,272 1,273
MMLU Anatomy 135
MMLU Clinical Knowledge 265
MMLU College Medicine 173
MMLU Medical Genetics 100
MMLU Professional Medicine 272
MMLU College Biology 144
Legal
BillSum 18,949 256
CAIL2018 (Chinese) 110,905 14,147 27,484
MMLU International Law 121
MMLU Jurisprudence 108
MMLU Professional Law 1,534
LexGLUE CaseHOLD 45,000 3,900 3,600
LexGLUE ECtHR (Task A) 6,838 802 808
General
ARC (Easy) 2,251 570 2,376
ARC (Challenge) 1,119 299 1,172
BoolQ 9,427 3,270
HellaSwag 39,905 10,042
LAMBADA (Standard) 4,869 5,153
OpenBookQA 4,957 500 500
PIQA 16,113 1,838
RTE 2,490 277
WinoGrande 40,398 1,267
StoryCloze 360 1,511
Table 3: Number of examples in each evaluation task.
Appendix D Infrastructure
We implement all experiments using PyTorch (Paszke et al., 2019) with the model implementations from Hugging Face Transformers (Wolf et al., 2020). We additionally use Hugging Face Datasets (Lhoest et al., 2021) for all dataset manipulation, including for tasks implemented via the EleutherAI LM Evaluation Harness (Gao et al., 2024). Finally, we conduct all experiments using a single NVIDIA A100 (SXM4 80GB) GPU.
Task Input Answer
CaseHOLD Given the following excerpt from a United States court opinion: Holding B
Citing Text:
… Warner-Lambert Co., 427 Mass. at 49 (“ [Confidential and proprietary business information may be entitled to protection, even if such information cannot claim trade secret protection”); see, e.g., Augat, Inc., 409 Mass. at 173 (<HOLDING>). “Matters of public knowledge or of general…
Given the following excerpt from a United States court opinion:
Which one of the following options should replace the <HOLDING> placeholder?
Holdings:
A. Recognizing that even if a plaintiff claims certain information constitutes trade secrets its claim may not depend on that determination.
B. Holding that included among trade secrets employee may not appropriate from employer is certain information such as lists of customers.
C. …
ECtHR Task A Given the following facts from a European Court of Human Rights (ECtHR) case: Choice K
Articles:
2. On 8 May 1996 the applicant was arrested in New York (USA) and placed in detention on the basis of a extradition request from the authorities of the Netherlands Antilles where …
3. In a document dated 18 June 1996 bearing the applicant’s …
Which article of the European Convention on Human Rights (ECHR) has been violated?
A. Article 2
B. Article 3
C. Article 5
K. None of the above
Jurisprudence The following are multiple choice questions (with answers) about jurisprudence: Choice B
Question:
Which statement best explains the purpose of Hart’s distinction between ’being obliged’ and ’having an obligation’?
Choices:
A. It demonstrates the difference between the internal and the external aspect of a rule.
B. It refutes the natural lawyer’ …
C. …
International / Professional Law The following are multiple choice questions (with answers) about professional / international law: Choice B
Statement:
One afternoon, a pilot was flying a small airplane when it suddenly ran out of gas. As he was coming in for an emergency landing … The attorney’s testimony is:
Choices:
A. admissible, because the…
B. inadmissible because …
C. …
Table 4: Representative task examples from the legal benchmark. Italicized text denotes the prompt.
Task Input Answer
MMLU Tasks: Clinical Knowledge; College Medicine; College Biology; Professional Medicine; Anatomy; Medical Genetics. The following are multiple choice questions (with answers) about <TASK>. Choice A
Question:
What size of cannula would you use in a patient who needed a rapid blood transfusion (as of 2020 medical knowledge)?
Choices:
A. 18 gauge
B. 20 gauge
C. …
PubMedQA Abstract Choice A - Yes
To evaluate the degree to which histologic chorioamnionitis, a frequent finding in placentas submitted for histopathologic evaluation, correlates with clinical indicators of infection in the mother. A retrospective review was performed on 52 cases with a histologic diagnosis of acute chorioamnionitis from 2,051 …
Question
Does histologic chorioamnionitis correspond to clinical chorioamnionitis?
Choices:
A. Yes
B. No
C. Maybe
MedMCQA Question: Choice D
All of the following are surgical options for morbid obesity except -:
Choices:
A. Adjustable gastric banding
B. Biliopancreatic diversion
C. Duodenal Switch
D. …
MedQA Question: Choice A
A 5-year-old girl is brought to the clinic by her mother for excessive hair growth. Her mother reports that for the past 2 months she has noticed … studies demonstrates an elevated level of estrogen. What is the most likely diagnosis?
Choices:
A. Granulosa cell tumor
B. Idiopathic precocious puberty
C. …
Table 5: Representative task examples from the biomedical benchmark. Italicized text denotes the prompt.
Task Input Answer
CAIL2018 根据以下法律案件的事实: Choice A
<FACT>
请问中华人民共和国刑法中的哪一条适用于本案?
A. 第 <ARTICLE 1> 条
B. 第 <ARTICLE 2> 条
C. 第 <ARTICLE 3> 条
D. 第 <ARTICLE 4> 条
答案:
CMB 问题: <QUESTION> Choice A
A. <OPTION 1>
B. <OPTION 2>
C. <OPTION 3>
D. <OPTION 4>
E. <OPTION 5>
答案:
Table 6: Example format of tasks in Chinese-language evaluation (CAIL2018 for legal and CMB for biomedical).
Appendix E Complete Results
Table 7 presents the full benchmark results (accuracy and perplexity) corresponding to Figure 3. For D-Pruner, we present results using iterative blocking and a loss regularization of 0.001. Finally, in Table 10 we further decompose benchmark results into their constituent tasks for completeness.
Model Method Target General Legal Biomedical
Domain Accuracy Perplexity Accuracy Perplexity Accuracy Perplexity
Llama 3.2 3B - - 61.60.0 9.30.0 45.50.0 5.20.0 53.30.0 9.40.0
Magnitude - 44.20.0 50.70.0 17.40.0 24.90.0 30.10.0 62.20.0
Wanda - 54.40.2 15.40.1 26.00.2 8.90.0 35.50.1 15.50.0
SparseGPT - 57.50.4 13.80.1 28.60.7 8.20.1 38.21.1 14.70.1
D-Pruner Legal 52.20.2 16.00.1 27.80.6 8.20.0 - -
Biomedical 52.10.1 19.00.2 - - 39.00.6 16.30.1
SparseGPT + MixCal Legal 56.80.3 13.60.0 36.50.7 7.70.2 - -
Biomedical 56.80.3 13.30.0 - - 44.70.4 13.50.1
Llama 3.1 8B - - 67.80.0 7.30.0 55.10.0 4.20.0 60.80.0 8.00.0
Magnitude - 47.60.0 57.70.0 20.90.0 65.70.0 35.40.0 55.70.0
Wanda - 60.60.1 11.70.0 35.00.3 6.80.0 42.40.2 11.90.0
SparseGPT - 62.60.2 10.70.1 40.10.6 6.30.1 45.81.4 11.50.1
D-Pruner Legal 57.20.3 12.60.2 42.51.8 6.70.0 - -
Biomedical 59.10.2 13.50.1 - - 49.20.4 14.70.3
SparseGPT + MixCal Legal 62.20.2 10.70.0 48.80.8 6.00.1 - -
Biomedical 63.40.1 10.50.0 - - 52.60.4 10.70.0
Gemma 2 2B - - 63.30.0 13.10.0 27.10.0 6.20.0 44.20.0 15.00.0
Magnitude - 48.20.0 172.50.0 21.70.0 34.70.0 26.80.0 825.80.0
Wanda - 54.20.2 25.00.2 20.11.5 10.70.0 29.60.7 33.00.5
SparseGPT - 56.60.2 20.90.4 20.81.4 9.20.1 30.41.3 26.60.3
SparseGPT + MixCal Legal 57.70.4 18.90.1 25.00.9 8.30.0 - -
Biomedical 58.10.2 19.10.1 - - 39.40.2 21.50.2
Gemma 2 9B - - 70.20.0 10.60.0 56.20.0 4.80.0 62.80.0 12.00.0
Magnitude - 60.60.0 33.50.0 33.70.0 11.00.0 47.80.0 57.40.0
Wanda - 63.50.2 16.60.1 41.81.2 6.70.0 48.80.4 20.80.2
SparseGPT - 65.80.2 15.20.2 43.70.8 6.40.0 52.80.5 19.30.2
SparseGPT + MixCal Legal 66.10.1 14.60.0 52.80.3 5.90.0 - -
Biomedical 66.50.1 14.50.1 - - 57.60.4 16.80.1
Mistral NeMo 12B - - 69.40.0 7.10.0 57.50.0 4.30.0 58.40.0 7.60.0
[…]
Wanda - 63.20.2 10.30.0 42.30.8 6.00.0 45.30.3 10.60.0
SparseGPT - 65.40.4 9.40.0 44.51.8 5.60.0 46.40.6 9.80.0
SparseGPT + MixCal Legal 65.30.2 9.50.0 51.11.1 5.50.0 - -
Biomedical 66.20.1 9.40.0 - - 53.70.4 9.60.0
Table 7: Average performance for the general and domain-specific benchmarks when pruning with 50% sparsity. Standard deviations are shown as subscripts. Bold values denote the best performing method for each model. For reference, the top row of each model shows the original model performance (i.e. prior to compression).
Model Method Target General Legal Biomedical
Domain Accuracy Perplexity Accuracy Perplexity Accuracy Perplexity
Llama 3.2 3B - - 61.60.0 9.30.0 45.50.0 5.20.0 53.30.0 9.40.0
SparseGPT - 49.60.2 22.60.2 21.71.0 14.60.2 27.30.4 25.70.2
SparseGPT + MixCal Legal 49.10.1 23.30.1 29.91.1 12.70.2 - -
Biomedical 50.10.1 22.80.1 - - 37.60.5 23.00.3
SparseGPT + GPTQ-M - 47.60.4 24.50.4 20.70.7 16.80.5 27.71.0 28.60.2
SparseGPT + GPTQ-M + MixCal Legal 46.00.6 25.20.3 28.92.5 14.00.3 - -
Biomedical 46.10.4 24.90.3 - - 35.60.4 25.30.4
Llama 3.1 8B - - 67.80.0 7.30.0 55.10.0 4.20.0 60.80.0 8.00.0
SparseGPT - 54.50.2 17.90.1 22.11.2 10.70.2 34.80.3 19.30.2
SparseGPT + MixCal Legal 53.60.1 18.20.2 41.50.8 9.60.1 - -
Biomedical 54.90.5 17.60.1 - - 42.40.7 17.00.2
SparseGPT + GPTQ-M - 54.20.3 19.30.6 20.71.4 12.00.2 33.90.5 20.90.2
SparseGPT + GPTQ-M + MixCal Legal 51.90.1 19.40.2 39.50.6 10.80.2 - -
Biomedical 52.80.3 19.00.2 - - 39.61.9 18.20.3
Gemma 2 2B - - 63.30.0 13.10.0 27.10.0 6.20.0 44.20.0 15.00.0
SparseGPT - 47.10.6 40.01.7 16.30.0 17.10.7 30.40.8 57.92.4
SparseGPT + MixCal Legal 48.20.3 34.40.4 21.40.5 13.00.1 - -
Biomedical 49.20.4 36.20.4 - - 32.00.8 40.70.8
SparseGPT + GPTQ-M - 45.80.5 45.22.0 16.30.1 19.20.8 30.42.1 68.35.8
SparseGPT + GPTQ-M + MixCal Legal 46.90.2 37.70.4 19.40.7 14.50.2 - -
Biomedical 47.60.2 40.10.5 - - 32.50.3 45.01.7
Gemma 2 9B - - 70.20.0 10.60.0 56.20.0 4.80.0 62.80.0 12.00.0
SparseGPT - 56.91.0 22.20.5 32.12.9 9.60.1 38.90.9 30.00.7
SparseGPT + MixCal Legal 57.60.2 21.10.1 44.61.1 7.80.0 - -
Biomedical 59.30.4 21.70.1 - - 49.20.6 23.60.1
SparseGPT + GPTQ-M - 56.10.4 24.30.4 27.52.8 10.40.1 36.51.4 33.80.6
SparseGPT + GPTQ-M + MixCal Legal 56.40.5 22.90.1 44.21.5 8.20.1 - -
Biomedical 58.50.6 23.50.1 - - 48.91.0 26.10.2
Mistral NeMo 12B - - 69.40.0 7.10.0 57.50.0 4.30.0 58.40.0 7.60.0
SparseGPT - 57.40.4 15.60.1 24.91.4 9.10.1 33.11.2 16.20.1
SparseGPT + MixCal Legal 56.60.2 16.30.1 38.00.5 8.20.1 - -
Biomedical 57.40.1 16.00.1 - - 43.70.3 15.10.1
SparseGPT + GPTQ-M - 56.10.2 17.30.3 22.21.5 10.20.2 32.02.3 17.90.3
SparseGPT + GPTQ-M + MixCal Legal 54.90.3 18.40.2 35.22.2 9.20.2 - -
Biomedical 56.00.1 18.10.5 - - 41.80.6 16.60.3
Table 8: Average performance for the general and domain-specific benchmarks when pruning with 2:4 sparsity. Standard deviations are shown as subscripts. Rows using the GPTQ-M method additionally employ 4-bit quantization. Bold values denote the best performing method per model for (1) pruning, and (2) joint pruning and quantization. For reference, the top row of each model shows the original model performance (i.e. prior to compression).
Group Size Model Target $\lambda_{g}$ General Legal Biomedical
Domain Accuracy Perplexity Accuracy Perplexity Accuracy Perplexity
None Llama 3.2 3B Legal 0.1 52.40.2 16.00.0 27.30.4 8.20.0 - -
Biomedical 0.1 50.90.6 18.00.0 - - 38.71.3 16.10.1
Legal 0.01 52.20.4 16.00.0 27.30.5 8.20.0 - -
Biomedical 0.01 51.00.6 17.90.1 - - 39.20.9 16.20.1
Legal 0.001 52.20.2 16.00.1 27.80.6 8.20.0 - -
Biomedical 0.001 50.90.5 17.80.1 - - 38.91.1 16.10.1
Llama 3.1 8B Legal 0.1 57.00.3 12.50.1 42.22.5 6.60.0 - -
Biomedical 0.1 57.60.7 14.00.4 - - 47.40.3 15.00.5
Legal 0.01 57.20.3 12.60.2 42.51.8 6.70.0 - -
Biomedical 0.01 58.80.2 13.30.1 - - 48.80.8 15.10.4
Legal 0.001 57.00.4 12.60.2 41.72.2 6.70.1 - -
Biomedical 0.001 57.71.2 14.21.0 - - 47.51.8 15.50.5
128 Llama 3.2 3B Legal 0.1 52.00.2 16.70.1 27.30.7 8.20.0 - -
Biomedical 0.1 52.00.2 18.90.2 - - 39.20.7 16.30.1
Legal 0.01 52.00.3 16.60.1 27.40.6 8.20.0 - -
Biomedical 0.01 52.10.1 19.00.2 - - 39.00.6 16.30.1
Legal 0.001 52.10.1 16.60.1 27.20.6 8.20.0 - -
Biomedical 0.001 52.00.3 18.90.2 - - 39.20.5 16.20.1
Llama 3.1 8B Legal 0.1 57.30.3 12.90.1 42.02.0 6.60.0 - -
Biomedical 0.1 58.10.8 14.20.4 - - 47.80.8 14.70.5
Legal 0.01 57.40.3 12.90.2 42.11.3 6.60.0 - -
Biomedical 0.01 59.10.2 13.50.1 - - 49.20.4 14.70.3
Legal 0.001 57.10.3 13.00.2 40.62.6 6.70.1 - -
Biomedical 0.001 58.11.2 14.30.9 - - 47.81.7 15.10.6
Table 9: Average D-Pruner performance for general and domain-specific benchmarks. We vary the group size (i.e. iterative blocking) and regularization ($\lambda_{g}$) hyperparameters. Standard deviations are shown as subscripts. Bold values denote the hyperparameter combination with the highest accuracy for each model and target domain.
Model Method Biomedical Legal
MedMCQA MedQA (4) PubMedQA Bio. MMLU CaseHOLD ECtHR Legal MMLU
Llama 3.2 3B - 49.50.0 51.50.0 72.80.0 61.10.0 42.70.0 49.60.0 44.20.0
Magnitude 28.80.0 28.80.0 50.60.0 27.50.0 20.60.0 4.50.0 27.20.0
Wanda 30.40.2 35.60.6 63.60.5 41.80.3 28.90.7 13.70.7 35.40.4
SparseGPT 33.61.9 35.41.4 69.51.2 45.11.2 32.90.6 15.81.6 37.10.5
D-Pruner 35.30.8 39.20.3 66.81.2 41.50.6 22.91.3 25.60.1 33.10.7
SparseGPT + MixCal 41.10.4 41.90.8 70.80.8 51.31.1 40.21.3 28.33.5 40.80.8
Llama 3.1 8B - 56.40.0 60.10.0 75.80.0 71.70.0 51.90.0 60.90.0 52.40.0
Magnitude 32.40.0 34.20.0 59.60.0 37.00.0 20.40.0 10.00.0 32.20.0
Wanda 38.50.3 38.00.9 66.80.5 51.30.6 29.40.7 35.90.3 39.70.3
SparseGPT 41.01.8 43.21.0 70.71.0 55.71.1 38.51.5 38.01.5 43.90.5
D-Pruner 43.41.5 45.12.1 69.92.8 57.52.0 37.12.4 41.46.6 43.40.8
SparseGPT + MixCal 48.30.5 49.20.8 72.30.9 63.40.6 45.60.5 54.21.4 46.61.0
Gemma 2 2B - 40.90.0 35.30.0 74.00.0 53.80.0 32.50.0 8.50.0 40.40.0
Magnitude 22.50.0 24.10.0 56.40.0 32.90.0 21.20.0 17.90.0 25.90.0
Wanda 25.80.9 24.40.9 57.40.6 37.61.0 19.80.1 12.24.8 28.30.6
SparseGPT 26.02.1 27.72.0 60.41.9 36.91.2 24.02.1 6.22.4 32.31.7
SparseGPT + MixCal 36.60.2 33.41.1 67.01.1 45.80.6 32.22.2 5.11.1 37.80.3
Gemma 2 9B - 57.90.0 60.50.0 78.60.0 77.20.0 51.70.0 60.10.0 56.80.0
Magnitude 43.70.0 46.30.0 71.80.0 54.50.0 37.10.0 19.90.0 44.10.0
Wanda 45.00.5 44.71.5 71.40.7 57.70.7 43.61.6 36.12.1 45.81.0
SparseGPT 48.00.5 48.60.9 74.41.1 66.10.4 45.81.5 36.20.9 49.20.9
SparseGPT + MixCal 53.00.4 54.11.1 76.40.9 70.50.2 49.80.1 55.00.4 53.60.5
Mistral NeMo 12B - 52.40.0 60.60.0 74.40.0 71.90.0 55.80.0 62.00.0 54.80.0
Magnitude 30.00.0 33.00.0 67.00.0 42.30.0 25.80.0 9.80.0 37.90.0
Wanda 39.70.4 45.50.3 60.60.1 59.70.4 44.00.5 37.42.8 45.40.5
SparseGPT 40.00.5 46.50.9 72.00.7 59.21.5 40.53.1 49.33.4 43.80.9
SparseGPT + MixCal 47.70.7 55.30.4 73.20.5 65.20.7 48.32.8 55.11.1 49.90.5
Table 10: Average performance across calibration sets for tasks from the biomedical and legal domains. Standard deviations are shown as subscripts. Bold values denote the best performing method for each task.
Model Method Target Domain General Legal Biomedical
Llama 3.2 3B SparseGPT - 13.80.1 8.20.1 14.70.1
+ In-domain data Legal 13.70.0 7.60.1 -
Biomedical 13.50.0 - 13.70.1
+ MixCal (Ours) Legal 13.60.0 7.70.2 -
Biomedical 13.30.0 - 13.50.1
Llama 3.1 8B SparseGPT - 10.70.1 6.30.1 11.50.1
+ In-domain data Legal 10.80.0 6.00.0 -
Biomedical 10.70.0 - 10.90.0
+ MixCal (Ours) Legal 10.70.0 6.00.1 -
Biomedical 10.50.0 - 10.70.0
Gemma 2 2B SparseGPT - 20.82.0 9.30.8 24.31.8
+ In-domain data Legal 19.90.1 8.30.0 -
Biomedical 20.50.1 - 23.50.2
+ MixCal (Ours) Legal 18.90.1 8.30.0 -
Biomedical 19.10.1 - 21.50.2
Gemma 2 9B SparseGPT - 15.20.2 6.40.0 19.30.2
+ In-domain data Legal 14.70.0 5.90.0 -
Biomedical 14.90.0 - 17.50.1
+ MixCal (Ours) Legal 14.60.0 5.90.0 -
Biomedical 14.50.1 - 16.80.1
Mistral NeMo 12B SparseGPT - 9.40.0 5.60.0 9.80.0
+ In-domain data Legal 9.50.0 5.50.0 -
Biomedical 9.50.0 - 9.70.0
+ MixCal (Ours) Legal 9.50.0 5.50.0 -
Biomedical 9.40.0 - 9.60.0
Table 11: Average performance when pruning to 50% unstructured sparsity with (1) in-domain data, and (2) MixCal. Standard deviations are shown as subscripts. Bold values denote the best performing method for each model.